OpenStudy (anonymous):

What is the solution to the equation? 4x – 6 = –3x + 8 @itsbribro

Vocaloid (vocaloid):

add 3x to both sides

OpenStudy (anonymous):

a.2/7 b.7/2 c.14 d.2

OpenStudy (itsbribro):

d which is 2 :)

OpenStudy (anonymous):

\[4x-6=-3x+8\] -3x +3x 7x-6 = 8 +6=+6 7x=14 ___ ___ 7 7 x=2